What is the pH of the solution containing 0.20 M NH3 and 0.15M NH4Cl?


1
Expert's answer
2022-05-30T17:02:05-0400

Kb (NH3)= 1.8*10-5 and pKb = 4.74


pOH = pKb - log ([NH4Cl] / [NH3] )

pOH = 4.74 - log (0.15/0.2)

pOH = 4.74 - log 0.75

pOH = 4.74 + 0.12

pOH = 4.86


pH = 14 - pOH

pH = 14.00 - 4.86

pH = 9.14